ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 41A17, 41A25, 41A35, 41A65, 42A10 Bustamante, J., and Jimenez, M. A., BUAP Puebla, Mexico The Degree of Best Approximation in the Lipschitz Norm by Trigonometric Polynomials We introduce a lipschitzian modulus of smoothness and obtain Jackson and Bernstein type theorems for trigonometric polynomial approximation in Lipschitz norm. A comparison with the usual modulus of smoothness is included.